### v1

# HJRNO: Hamilton-Jacobi Reachability with Neural Operators

Abstract
Ensuring the safety of autonomous systems under uncertainty is a critical challenge. Hamilton-Jacobi reachability (HJR) analysis is a widely used method for guaranteeing safety under worst-case disturbances. Traditional HJR methods provide safety guarantees but suffer from the curse of dimensionality, limiting their scalability to high-dimensional systems or varying environmental conditions. In this work, we propose HJRNO, a neural operator-based framework for solving backward reachable tubes (BRTs) efficiently and accurately. By leveraging the Fourier Neural Operator (FNO), HJRNO learns a mapping between value functions, enabling fast inference with strong generalization across different obstacle shapes, system configurations, and hyperparameters. We demonstrate that HJRNO achieves low error on random obstacle scenarios and generalizes effectively across varying system dynamics. These results suggest that HJRNO offers a promising foundation model approach for scalable, real-time safety analysis in autonomous systems.
1 Introduction
Autonomous systems are playing an increasingly significant role in daily life, drawing growing attention from both industry and academia. Despite their versatile capabilities and the substantial assistance they provide, ensuring their safety remains a critical concern. Many autonomous systems are exposed to unpredictable disturbances, such as varying weather conditions, which can impact their reliability and performance. Hamilton-Jacobi (HJ) reachability analysis has emerged as a powerful tool for providing provable safety guarantees under worst-case disturbances, addressing both system configurations and control strategies [1, 6].
HJ reachability analysis achieves this by computing the backward reachable tube (BRT), which represents the set of states that the system must avoid to maintain safety. In addition to identifying unsafe regions, the BRT offers a quantitative measure of how close the current configuration is to unsafe states. Furthermore, the process of computing the BRT naturally yields an optimal control strategy that directs the system away from the unsafe region.
The computation of the BRT involves solving the Hamilton-Jacobi (HJ) partial differential equation (PDE) over the state space, typically discretized using a grid-based dynamic programming approach. This traditional method can produce highly accurate results when a high-resolution grid is used. However, the “curse of dimensionality” poses a significant challenge, as the number of grid points grows exponentially with the dimension of the system. Moreover, solving the HJ-PDE requires cross-dimensional computations, further compounding the computational burden.
Several research directions have been explored to accelerate the computation of the BRT. One approach involves computing an approximate BRT by decomposing the system into smaller subsystems, thereby avoiding cross-dimensional interactions [5, 4]. Another strategy simplifies the system dynamics, such as by linearizing nonlinear systems [12]. While these approximations can significantly reduce computational costs, they often lead to inaccuracies in the resulting BRT and typically require additional system-specific analysis to ensure safety.
Recently, deep learning-based approaches have been proposed to efficiently compute the BRT. The physics-informed machine learning framework [2, 3, 13, 14] introduces neural PDE solvers that leverage neural networks to compute the BRT. However, these frameworks only learn the solution function for a single instance rather than the underlying functional mapping. Consequently, when the input function changes, a new neural PDE solver must be trained from scratch.
In this work, we propose solving the BRT using neural operators [9, 11, 10], which learn mappings between infinite-dimensional function spaces. This approach offers two advantages: (1) the neural operator model needs to be trained only once and can subsequently generalize across a broad range of problem settings; (2) inference time is orders of magnitude faster than traditional BRT solvers.
Our main contributions are as follows:
  1. 1.
    To the best of our knowledge, this is the first work to apply neural operators to Hamilton-Jacobi reachability (HJR) problems.
  2. 2.
    By learning a neural operator, we achieve near-instantaneous inference ( $10^{-3}$ seconds), eliminating the need to solve the HJ-PDE for each different problem setting, such as varying obstacle shapes or system hyperparameters.
  3. 3.
    We demonstrate that our method generalizes well across randomly generated obstacle shapes and varying system hyperparameters, indicating strong potential for applying neural operators to BRT problems.
2 Problem Setting
We consider the dynamics of an autonomous system described by
$\dot{x}=f(x,u,d),$ (1a)
$x(0)=x_{0},$ (1b)
$x\in\mathbb{R}^{n},\;u\in\mathcal{U},\;d\in\mathcal{D}$ (1c)
where $x$ denotes the system state, $u$ is the control input, and $d$ represents an external disturbance. The sets $\mathcal{U}$ and $\mathcal{D}$ denote the sets of measurable control and disturbance functions, respectively. We define the system trajectory starting from initial state $x$, subject to control $u(\cdot)$ and disturbance $d(\cdot)$ over the time interval $[t,\tau]$, as $\zeta(x,u,d,t,\tau)$. The unsafe set is defined as =$\mathcal{L}=\{x\;:\;l(x)\leq 0\}$. The backward reachable tube (BRT) is the set of all initial states from which, despite applying optimal control strategies aimed at avoiding $\mathcal{L}$, there exists a disturbance strategy that forces the system into the unsafe set within the time horizon $[t,\tau]$:
$\mathcal{R}(t)=\{x:\forall u(\cdot),\exists d(\cdot),\exists\tau\in[t,T],\zeta(x,u,d,t,\tau)\in\mathcal{L}\}$ (2)
Since the backward reachable tube (BRT) contains unsafe states, the objective in system design is to steer trajectories away from the BRT. To compute the BRT, we first define an objective function that measures the minimum value of $l$ along a trajectory:
$J(x,t,u,d)=\underset{\tau\in[t,T]}{\mathrm{min}}l(\zeta(x,u,d,t,\tau))$ (3)
The control strategy seeks to maximize this objective, effectively pushing the system away from unsafe states, while the disturbance seeks to minimize it. This leads to the definition of the value function:
[…]
$V(x,T)$ $=l(x)$ (4b)
The value function $V(\cdot)$ is the viscosity solution of the Hamilton-Jacobi-Isaacs (HJI) partial differential equation (PDE) [8, 7]:
$\min\left\{D_{t}V(x,t)+H(x,t),\;l(x)-V(x,t)\right\}=0$ (5)
where the Hamiltonian $H(x,t)$ is defined as:
[…]
As an illustrative example, we model an autonomous vehicle using a generalized Dubins car model. Unlike the classical Dubins car, this formulation allows the vehicle’s speed to vary and adopts a more realistic heading control mechanism based on curvature, which is standard for nonholonomic vehicles. Specifically, rather than directly controlling the turning rate, we apply a curvature control scaled by the velocity. In the following, we refer to this Dubins-like vehicle with acceleration and curvature control as the dynamic Dubins car. The system dynamics are given by:
$\dot{x}_{1}=v\cos(\theta)+d_{1},$ (7a)
$\dot{x}_{2}=v\sin(\theta)+d_{2},$ (7b)
$\dot{v}=u_{1},$ (7c)
$\dot{\theta}=v\,u_{2}$ (7d)
where the states are the position components $(x_{1},x_{2})$, the velocity $v$, and the heading angle $\theta$. The control inputs are the acceleration $u_{1}$ and the curvature $u_{2}$, while $d_{1}$ and $d_{2}$ represent disturbances acting on the $x_{1}$ and $x_{2}$ positions, respectively.
To define an unsafe set, consider the task of avoiding a tree obstacle. The unsafe set can be modeled as:
$l(x)=(x_{1}-x_{\mathrm{tree}})^{2}+(x_{2}-y_{\mathrm{tree}})^{2}-r$ (8)
where $(x_{\mathrm{tree}},y_{\mathrm{tree}})$ denotes the center of the tree, and $r$ represents its effective radius. The BRT can then be computed by solving the HJI-PDE (5). Although this example assumes the obstacle (tree) has a circular shape, in the Results section, we demonstrate that our proposed method can handle obstacles of arbitrary shapes effectively.
3 Method
3.1 Neural Operator
Operator learning focuses on learning mappings between coefficient functions (inputs) and solution functions (outputs). A straightforward approach would be to discretize these functions onto grids, allowing the use of conventional neural network architectures such as multilayer perceptrons (MLPs), convolutional neural networks (CNNs), or vision transformers (ViTs) for learning. However, this strategy inherently ties the model to the specific discretization used during training, limiting its ability to generalize across different resolutions or grid structures.
In contrast, neural operators are designed to directly learn mappings between function spaces without relying on fixed discretizations. As a result, neural operators can naturally adapt to new discretizations and varying resolutions without the need for retraining. Moreover, traditional neural networks often require significant architectural modifications to maintain a consistent level of accuracy across different discretizations. Neural operators are the only known models that offer a theoretical guarantee of both universal approximation and invariance to discretization [9].
The Fourier transform simplifies differentiation by converting it into multiplication in the Fourier domain, making it a popular tool for solving PDEs. Building on this principle, the Fourier Neural Operator (FNO) [11] has emerged as one of the leading architectures for learning operators between function spaces. The FNO constructs the mapping using a sequence of $L$ integral kernel operator layers:
$s=\mathcal{M}_{\theta}(a):=\mathcal{Q}\circ\sigma(\mathcal{W}_{L}+\mathcal{K}_{L})\circ\cdots\circ\sigma(\mathcal{W}_{1}+\mathcal{K}_{1})\circ\mathcal{P}(a)$ (9)
where $a(\cdot)$ and $s(\cdot)$ are the input and output functions, $\sigma$ is an activation function, $\mathcal{P}$ and $\mathcal{Q}$ are shallow fully connected networks for projecting to and from higher-dimensional spaces, and $\mathcal{W}_{i}$ are learned linear maps. Each $\mathcal{K}_{i}$ represents a Fourier integral operator to the intermediate layer output $v_{i}$:
$(\mathcal{K}_{i}v_{i})(x)=\mathcal{F}^{-1}\big{(}R_{i}\cdot(\mathcal{F}v_{i})\big{)}(x)$ (10)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier and inverse Fourier transforms, respectively, and $R_{i}$ is a learnable weight matrix that parameterizes the kernel function in the Fourier space. By utilizing frequency-domain representations, FNO efficiently capture both local and global behaviors of complex functions.
3.2 HJRNO
We apply the Fourier Neural Operator (FNO) [11] to learn the functional map for backward reachable tubes (BRT):
$a(x)$ $:=V(x,\,t=T)$ (11a)
$s(x)$ $:=V(x,\,t=0)$ (11b)
To construct the training dataset, we first compute the solutions using traditional dynamic programming solvers. We then train the FNO (9) in a purely data-driven manner to learn the mapping between these functions. Although intermediate time steps are not explicitly available, we set $T$ sufficiently large so that $V(x,\,t=0)$ stabilizes and does not change with further increases in $T$. This setup reflects the most common practical use case.
Once trained, our method, which we call HJRNO (Hamilton-Jacobi Reachability with Neural Operators), achieves an inference time of approximately $\sim 10^{-3}$ seconds—orders of magnitude faster than traditional solvers, which typically require several seconds to hours per evaluation.
Additionally, to account for hyperparameters $h$ (e.g., the maximum acceleration limit in the dynamic Dubins car model from the Problem Setting section), we extend the input by appending $h$ as a constant feature across all spatial locations:
$a(x)$ $:=V(x,\,t=T,h)$ (12a)
$s(x)$ $:=V(x,\,t=0,h)$ (12b)
Leveraging the universal approximation and discretization invariance properties of FNO, HJRNO can predict solutions for a wide range of $V(x,\,t=T,h)$ instances. As we show in the Results section, HJRNO exhibits strong interpolation capability even with a limited training set, greatly reducing computational costs when handling varying obstacle shapes or system hyperparameters—scenarios where traditional BRT recomputation would be expensive. Finally, thanks to discretization invariance, HJRNO also enables zero-shot super-resolution: it can predict high-resolution solutions even when trained only on low-resolution datasets.
4 Results
Ground TruthRefer to captionRefer to captionRefer to captionPredictionRefer to captionRefer to captionRefer to captionError MapRefer to captionRefer to captionRefer to caption
Figure 1: Predicted BRTs for randomly selected test cases from the random obstacle shape experiment
We evaluate HJRNO on the dynamic Dubins car model introduced in the Problem Setting section. To simulate realistic and diverse environments, we generate random obstacles with smooth, arbitrary shapes, examples of which are shown in Fig.4. The model is trained on 100 such obstacle shapes. Across 1000 test cases, HJRNO achieves a low average error of 0.015. Results randomly selected from the test dataset are presented in Fig.1.
We further evaluate HJRNO’s ability to generalize across varying system hyperparameters, specifically the maximum acceleration limit and maximum curvature limit. In this experiment, we use a fixed circular obstacle shape. We sample a sparse training dataset by selecting 10 uniformly spaced values along each hyperparameter axis, yielding $10\times 10=100$ training points. The range of each hyperparameter is shown in Fig. 3.
Ground TruthRefer to captionRefer to captionRefer to captionRefer to captionRefer to captionPredictionRefer to captionRefer to captionRefer to captionRefer to captionRefer to caption
Figure 2: Predicted BRTs along the diagonal of the hyperparameter space
Figure 3: Sampling and testing strategy across varying system hyperparameters
To test generalization, we predict along the diagonal of the hyperparameter space. Fig. 2 illustrates the results, where predictions are made for random hyperparameter values along this diagonal corresponding to hyperparameter values not present in the training set.
Although we visualize only slices of the BRT over the $x_{1}$ (position $x$) and $x_{2}$ (position $y$) dimensions, all predictions in this section are performed over the full four-dimensional state space, including velocity $v$ and heading angle $\theta$. The results shown are obtained by randomly slicing along the $v$ and $\theta$ dimensions.
Figure 4: Random obstacle shapes
5 Discussion
The results demonstrate that HJRNO performs impressively across diverse tasks, suggesting strong potential as a foundation model for solving BRT problems.
While we presented results on a uniform grid, the method naturally extends to arbitrary geometries [10]. This means HJRNO can be directly applied to environments with non-square or even non-convex domains—for example, an autonomous agent operating within a warehouse with an irregular floor plan.
One limitation is the treatment of hyperparameters. When introducing additional hyperparameters, each value must be repeated across the input domain, increasing the size of the training dataset proportionally. This approach could become a computational bottleneck when scaling to high-dimensional hyperparameter spaces. A future research direction could be to design more efficient architectures that better handle hyperparameters, particularly those that are spatially constant, without requiring dataset duplication.